Circular No. 6202

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) Continuation to the ephemeris on IAUC 6191 (elements IAUC 6198):

1995 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Aug. 11 18^{h}31.^{m}67 −31^{o}34.^{′}6 6.207 6.995 138.^{o}2 5.^{o}5 10.4
21 18 26.26 −31 10.6 6.244 6.910 127.7 6.7 10.4
31 18 21.92 −30 44.4 6.302 6.824 117.3 7.6 10.3
Sept. 10 18 18.77 −30 16.9 6.373 6.738 107.1 8.2 10.3
20 18 16.86 −29 48.8 6.453 6.651  97.0 8.6 10.3
30 18 16.17 −29 20.6 6.537 6.564  87.1 8.8 10.2

Total magnitude estimates (visual unless otherwise stated): July 17.89 UT, 11.5 ± 0.5 (P. Riepe, B. Schrvter, H. Tomsik and S. Binnewies, Windhoek, Namibia; 0.135-m f/4.0 lens, Scotchchrome 400; prediscovery photograph); 18.92, 11.5 ± 0.5 (Riepe et al., prediscovery photograph); 31.09, 10.4 (J. Bortle, Stormville, NY, 0.41-m reflector); Aug. 1.11, 10.3 (Bortle); 2.18, 10.8 (A. Hale, Cloudcroft, NM, 0.41-m reflector).

The following near-perihelic ephemeris is provided. It must be stressed that the computation depends critically on the reality of the single 1993 Apr. 27 observation (IAUC 6198) and on the assumption that the comet has not been significantly in outburst (but see IAUC 6194). Taking the situation at face value, one can note that comet C/1995 O1 is in many respects similar to the great comet C/1811 F1 (early absolute brightness, perihelion distance, orbital inclination, 3000-year revolution period, placement beyond the sun at perihelion) and may perform as spectacularly.

1997 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Feb. 1 19^{h}40.^{m}35 +15^{o}18.^{′}8 2.012 1.373 37.^{o}8 26.^{o}1  0.9
11 20 07.44 +20 38.3 1.815 1.256 41.3 31.2  0.3
21 20 42.87 +27 08.3 1.627 1.148 44.2 36.9 −0.3
Mar. 3 21 32.05 +34 34.1 1.467 1.053 45.9 42.5 −0.9
13 22 42.51 +41 40.2 1.356 0.978 46.1 47.0 −1.4
23 0 16.14 +45 41.3 1.317 0.929 44.7 49.0 −1.7
Apr. 2 1 53.97 +44 21.1 1.358 0.913 42.3 47.4 −1.7
12 3 11.52 +38 54.5 1.465 0.932 39.0 42.6 −1.5
22 4 05.54 +32 08.6 1.615 0.983 35.2 36.2 −1.0
May 2 4 43.56 +25 38.4 1.782 1.059 31.3 29.6 −0.5
12 5 12.07 +19 51.2 1.953 1.155 27.5 23.8  0.1
22 5 34.94 +14 46.0 2.116 1.264 24.3 19.3  0.6
June 1 5 54.33 +10 13.7 2.266 1.381 22.2 16.1  1.2

1995 August 4 Brian G. Marsden